: In a hypothetical economy, without taxes, the consumption(C) at different levels of real GDP (Y) is
given here under:
Real GDP Consumption (Rs in Billion)
(Y) (C)
3,000 1,500
5,000 3,000
Determine the size of MPC (Marginal propensity to consume), MPS (Marginal propensity to save) and multiplier
1
Expert's answer
2020-09-16T20:51:10-0400

MPC is share of additional income that consumer spend on consumption.

MPC = "\frac{ΔC}{ΔY}"


"\frac{3000-1500}{5000-3000}."



="\frac{1500}{2000}"

MPC = 0.75


MPS

MPC+ MPS= 1 Since there is no tax for this case.


"0.75+MPS= 1"

MPS = "1 - 0.75"


MPS = 0.25.


Multiplier

denoted as k

k= "\frac{1}{1-mpc}"


k="\frac{1}{1-0.75}"


k ="\frac{1}{0.25}"

k= 4.